# Toward thermoelectric characterization of (nano)materials by in situ transmission electron microscopy

### v1

@@ Abstract
We explore the possibility to perform an in situ transmission electron microscopy (TEM) thermoelectric characterization of materials. A differential heating element on a custom in situ TEM microchip allows to generate a temperature gradient across the studied materials, which are simultaneously measured electrically. A thermovoltage was induced in all studied devices, whose sign corresponds to the sign of the Seebeck coefficient of the tested materials. The results indicate that in situ thermoelectric TEM studies can help to profoundly understand the interplay between the thermoelectric properties on one side and the structure/composition of the materials down to the atomic level on the other side, including grain boundaries, dopants or crystal defects. We propose an improved in situ TEM microchip design, which should facilitate a full quantitative measurement of the induced temperature gradient, the electrical and thermal conductivities, as well as the Seebeck coefficient.
@@
in-situ transmission electron microscopy, measurement on individual nanomaterial, sample preparation, electrical properties, focused ion beamkeywords:
in situ transmission electron microscopy, seebeck coefficient, thermoelectricity, nanomaterialSimon Hettler* Mohammad Furqan Andrés Sotelo Raul Arenal*
Dr. S. Hettler, M. Furqan, Dr. A. Sotelo, Dr. R. Arenal
Instituto de Nanociencia y Materiales de Aragón (INMA), CSIC-Universidad de Zaragoza, 50009 Zaragoza, Spain
Email Address: hettler@unizar.es, arenal@unizar.es
Dr. S. Hettler, M. Furqan, Dr. R. Arenal
Laboratorio de Microscopías Avanzadas (LMA), Universidad de Zaragoza, 50018 Zaragoza, Spain
Dr. R. Arenal
ARAID Foundation, 50018 Zaragoza, Spain
@@ 1 Introduction
Thermoelectric materials are an important building block for generation of green energy from heat and a huge range of different materials are being studied [1, 2, 3, 4]. In general, the thermoelectric effect is a complex phenomenon and the Seebeck coefficient strongly depends on crystallinity, presence of dopants, contaminants or impurities in the material [5, 6, 7]. In the development of thermoelectric bulk materials, different strategies are followed to improve the figure of merit, , including grain-boundary engineering [6] or doping [7]. Nanomaterials, nano-composites or nano-structured devices are also promising candidates for thermoelectric materials [8, 9, 5]. The miniaturization of devices, the use of nanomaterials and the complex interplay between the structure/composition of materials and their thermoelectric properties have led to numerous different approaches for thermoelectric characterization of (nano)materials [10, 11, 12, 13].
@@
Transmission electron microscopy (TEM), combined with related spectroscopic techniques, allows the in-depth analysis of structure and composition down to the atomic scale [14, 15]. Technological developments have improved the energy resolution in electron energy-loss spectroscopy (EELS) and allow the local study of phonons by scanning (S)TEM [16, 17], which are important for heat transport phenomena. Recently, in situ TEM studies have gained significant importance as they allow to study the evolution of materials under external stimuli. One branch of in situ experiments are electrical studies, where the influence of an applied voltage [18, 19, 20] or current [21] on a material is investigated. Here, we explore the possibility to measure the Seebeck coefficient and thermoelectric properties of materials, including individual nanomaterials, by in situ TEM. The combination of thermoelectric characterization and TEM is promising for gaining a deeper understanding of fundamental processes in thermoelectric materials as, for example, the role of grain boundaries.
@@ 2 Results and discussion
2.1 Device setup
The general concept to measure the Seebeck coefficient is to generate a temperature gradient along the material to be studied and simultaneously measure the generated electrical potential. To combine such a measurement with TEM investigations, a microchip that fits in an in situ TEM sample holder was designed and fabricated by microelectromechanical systems (MEMS) technologies. Chip fabrication is described in detail in the experimental section. Figure 1 shows a photograph, an optical microscopy image and a scanning electron microscopy (SEM) image of one of these chips, which consist of a differential heating device and two contact pads (10 nm Ti + 150 nm Pt) on top of a free-standing low-stress silicon nitride membrane (thickness 1 m). The differential heating device is located on one side of the membrane in order that an applied heating current () would create a temperature gradient between the contact pads, which are located in the center of the membrane. The polarity of the voltage applied to the device is indicated in Figure 1c.
We studied five different thermoelectric devices with the fabricated chips. Four devices were prepared from bulk samples: a standard p-doped (boron) silicon single crystal, single-crystalline molybdenum (source: sputter target, structural analysis in supplementary information (SI), section SI1) and a strontium-doped calcium cobaltite (CCO: Ca2.93Sr0.07Co4O9) misfit-layered compound (MLC). For the CCO material, two devices were fabricated with different crystal orientation as discussed below. The fifth device is a nanotube (NT) based on the MLC (LaS)1.14-TaS2 [22]. In order to study these materials, a piece of each has to be deposited between the two contact pads on the in situ TEM chip. As detailed in the experimental section, the focused ion beam (FIB) technique was used to prepare the devices. Two different approaches have been employed for bulk [23] and NT devices [24], respectively.
@@ 2.2 Reference materials and device characterization
Figure 2a shows a SEM image of the prepared p-doped Si device, where the material was cut into a cuboid with defined geometry in the central area above the hole in the membrane. The crystal orientation is 100 in vertical direction and approximately 032 in direction of the current flow between the contacts (SAED analysis in SI2). Energy-dispersive X-Ray spectroscopy (EDX) indicates a minor surface oxidation and Ga incorporation (SI3). Figure 2b depicts an I-V characterization of the device shown in (a) obtained at different applied heating currents between 0.1 (blue curve) and 5 mA (red curve).
With increasing applied to the differential heating device, the I-V curve shifts along the x-axis and the gradient slightly increases. The gradient change is linked to a slight decrease in average resistance from 890 to 775 due to an increase in temperature and a related improvement of the semiconducting Si and also likely due to a reduction of contact resistance. The shift can be attributed to the generation of a voltage induced by a temperature gradient along the sample. To measure this shift, a linear regression is performed on the I-V curves and the shift is calculated from and is marked by black circles in Figure 2b. The negative sign is due to the employed polarity (Figure 1c), as the temperature gradient is opposite to the applied voltage. The voltage offset is determined by the value of relative to (=0) to account for possible minor intrinsic currents. is shown in Figure 2c in dependence of for the Si device and a similar sample prepared from single-crystalline Mo (SI4). Both curves show a clear increase of the voltage offset with , which corresponds with the expected increase of the temperature gradient along the material. The voltage offset is positive, which agrees with the (positive) sign of the Seebeck coefficient of both materials [25, 26]. A quantitative assessment of the measured voltages is discussed below.
To check a possible contribution to the thermolectrical measurement by the chip itself, the raw in situ chip, without deposited material, was characterized in the same way. No measurable contribution was observed (SI5).
@@ 2.3 Devices made of misfit-layered compounds
Figure 3 shows the structural analysis of the three MLC devices. Two devices were prepared from the bulk CCO material with different crystal orientations with respect to the induced temperature gradient. Two SEM images of the first CCO device before and after FIB thinning for TEM studies are shown in Figure 3a. The preparation of this device is documented in SI6. The microstructure of the Sr-doped CCO bulk material was analyzed using a standard TEM lamella (SI7), which revealed that the material possesses a micro-crystalline structure with most grains exhibiting a similar direction of the c-axis. The c-axis corresponds to the direction of the applied pressure during the hot-pressing process, see experimental section. Grain sizes are small in c-direction (approximately 500 nm) and large in a-b directions (several m). A few grains and grain boundaries spanning horizontally across the device can be seen in the TEM image in Figure 3b. Figure 3c depicts a selected-area electron diffraction (SAED) pattern obtained from the thin part of the device. The SAED pattern confirms that the c-axis is perpendicular to the grain boundaries visible in Figure 3b, implying that for this device, the electrical current flow is predominantly in-plane and thus parallel to the layers of the MLC. This device is denoted as CCO in the following. The c-axis periodicity of the CCO material is measured to be 1.07 nm from the SAED pattern, which agrees with previously reported values on bulk CCO [27].
@@
Figure 3d-g details a second CCO device, denoted as CCO , whose preparation is shown in SI8. Figure 3d shows a SEM image of the device, which was prepared such that the current flow is roughly perpendicular to the layered structure (parallel to the c-axis) and that a single grain boundary falls in the thinned central part. Figure 3e shows a bright-field TEM image of this device after orientation of the left grain into a high-order zone axis (TEM stage at =23°, =23°). The image clearly reveals the presence of two grains with the left grain oriented in zone-axis, appearing darker compared to the right grain. Figure 3f and 3g show two SAED patterns obtained from the left and right part of the device as marked in Figure 3e, corresponding to the two grains. The patterns confirm the orientation of the left grain in a higher-order zone-axis. An analysis of the reflections indicates the presence of [001] (1.07 nm) and [130] (0.15 nm) reflections, corresponding to a [-310] zone-axis orientation. Note that due to the monoclinic structure of the CCO material, [001] and [130] are not perpendicular to each other. On the other hand, the right grain does not exhibit a higher order orientation of the crystal with respect to the electron beam at this particular stage tilt (Figure 3g). The analysis of the device tilted to an angle where the right grain is oriented in [010] zone-axis (TEM stage at =16°, =19°) is shown in SI9. As the 001 direction roughly corresponds to the tilt direction, the angle between the c-axis directions of the two grains is estimated by the difference in tilt, which roughly corresponds to 7°. An EDX spectrum of this CCO device is shown in SI3 and confirms the presence of all elements of the CCO device, including the strontium doping.
@@
The third MLC device (preparation in SI10) consists in a NT made of (LaS)1.14-TaS2. A SEM image of the device is shown in Figure 3h. The silicon nitride membrane utilized for the sample transfer to the chip was milled away in the central region in order that the NT forms the only conductive connection between the contact pads. A rather long Pt-C focused ion beam induced deposition (FIBID) was needed to contact the NT to the upper contact pad. A SAED pattern acquired from the central part of the NT (Figure 3i) reveals the complex structure of the MLC NT [22]: Reflections are seen from the two different subsystems LaS (orange dashed rings) and TaS2 (blue dashed rings) and the c-axis periodicity of the stack (pink arrow), which is perpendicular to the NT axis, as seen in Figure 3h. The tubular structure is recognized from the streaking reflections caused by the bending of the layers. The MLC stack formed of bright TaS2 layers and LaS layers is clearly seen in a high-angle annular dark field (HAADF) STEM image (Figure 3j). Two amorphous surface layers are visible in the image. The inner, brighter part is intrinsic to the NT and was observed already prior to the transfer (SI11). The outer, dark layer was generated during the transfer process and is mainly composed of carbon and Pt stemming from the precursor gas used for the FIBID procedure employed to contact the NT. SI3 shows an STEM-EDX spectrum of the MLC-NT device confirming the expected composition.

### v2

# Toward quantitative thermoelectric characterization of (nano)materials by in-situ transmission electron microscopy

@@ Abstract
We explore the possibility to perform an in-situ transmission electron microscopy (TEM) thermoelectric characterization of materials. A differential heating element on a custom in-situ TEM microchip allows to generate a temperature gradient across the studied materials, which are simultaneously measured electrically. A thermovoltage was induced in all studied devices, whose sign corresponds to the sign of the Seebeck coefficient of the tested materials. The results indicate that in-situ thermoelectric TEM studies can help to profoundly understand fundamental aspects of thermoelectricity, which is exemplary demonstrated by tracking the thermovoltage during in-situ crystallization of an amorphous Ge thin film. We propose an improved in-situ TEM microchip design, which should facilitate a full quantitative measurement of the induced temperature gradient, the electrical and thermal conductivities, as well as the Seebeck coefficient. The benefit of the in-situ approach is the possibility to directly correlate the thermoelectric properties with the structure and chemical composition of the entire studied device down to the atomic level, including grain boundaries, dopants or crystal defects, and to trace its dynamic evolution upon heating or during the application of electrical currents.
@@
in-situ transmission electron microscopy , Seebeck coefficient , thermoelectricity , nanomaterialThis is the accepted version of the article published open access in Ultramicroscopy:
https://doi.org/10.1016/j.ultramic.2024.114071
@@ 1 Introduction
Thermoelectric materials are an important building block for generation of green energy from heat and a huge range of different materials are being studied [1, 2, 3, 4]. In general, the thermoelectric effect is a complex phenomenon and the Seebeck coefficient strongly depends on crystallinity, presence of dopants, contaminants or impurities in the material [5, 6, 7]. In the development of thermoelectric bulk materials, different strategies are followed to improve the figure of merit, , including grain-boundary engineering [6] or doping [7]. Nanomaterials, nano-composites or nano-structured devices are also promising candidates for thermoelectric materials [8, 9, 5]. The miniaturization of devices, the use of nanomaterials and the complex interplay between the structure/composition of materials and their thermoelectric properties have led to numerous different approaches for electric and thermoelectric characterization of (nano)materials [10, 11, 12, 13].
@@
Transmission electron microscopy (TEM), combined with related spectroscopic techniques, allows the in-depth analysis of structure and composition down to the atomic scale [14, 15]. Technological developments have improved the energy resolution in electron energy-loss spectroscopy (EELS) and allow the local study of phonons by scanning (S)TEM [16, 17], which are important for heat transport phenomena. Recently, in-situ TEM studies have gained significant importance as they allow to study the evolution of materials under external stimuli. One branch of in-situ experiments are electrical studies, where the influence of an applied voltage [18, 19, 20] or current [21] on a material is investigated. Here, we explore the possibility to measure the Seebeck coefficient and thermoelectric properties of materials, including individual nanomaterials, by in-situ TEM. The combination of thermoelectric characterization and TEM is promising for gaining a deeper understanding of fundamental relationships in thermoelectric materials as, for example, the role of grain boundaries.
2 Materials and Methods
2.1 Device setup
The general concept to measure the Seebeck coefficient is to generate a temperature gradient along the material to be studied and simultaneously measure the generated electrical potential. To combine such a measurement with TEM investigations, a microchip that fits in an in-situ TEM sample holder was designed and fabricated by microelectromechanical systems (MEMS) technologies. The chip fabrication procedure was described elsewhere in detail [22]. Figure 1 shows a photograph, an optical microscopy image and a scanning electron microscopy (SEM) image of one of these chips, which consist of a differential heating device and two contact pads (10 nm Ti + 150 nm Pt) on top of a free-standing low-stress silicon nitride membrane (thickness 1 m). The differential heating element is located on one side of the membrane in order that an applied heating current () would create a temperature gradient along the specimen placed between the contact pads (Figure 1c), which are located in the center of the membrane. To measure the voltage induced in thermoelectric materials by a specific temperature gradient, I-V curves are acquired, which provide higher accuracy compared to a simple voltage measurement and give additional information on the resistivity. The polarity of the voltage applied to the specimen is indicated in Figure 1c.
2.2 Investigated materials
To prove that reliable thermoelectric measurements are possible with the proposed in-situ approach and to illustrate the advantage of the approach, we studied five materials with different thermoelectric properties by preparing devices using the fabricated chips. We chose p-doped (boron) single-crystalline silicon and a single-crystalline molybdenum (source: sputter target, structural analysis in supplementary information (SI), section SI1) as reference standards for a semiconducting and metallic material, respectively. As a third well-studied reference material, strontium-doped calcium cobaltite (CCO: Ca2.93Sr0.07Co4O9) misfit-layered compound (MLC) was used. The CCO bulk material was prepared as described in [7] by attrition milling (2h at 600 rpm) starting from SrCO3, CaCO3 and CoO. The mixture was calcinated at 850 °C during 1 h, cold uniaxially pressed into 25 mm diameter disks under 250 MPa, and subjected to hot-uniaxial pressing at 900 °C and 51 MPa for 1 h. For this CCO material, two devices were fabricated with the aim to study the relationship between different crystal directions of the asymmetric material and its thermoelectric properties, as discussed below.
An example for a material with negative Seebeck coefficient are nanotubes (NTs) based on the MLC (LaS)1.14-TaS2 [23, 1]. Finally, two devices were prepared based on amorphous Ge (aGe) and amorphous C (aC) thin films to perform dynamic in-situ experiments (see section 3.5). aGe and aC/aGe/aC thin films were prepared by electron-beam evaporation (Auto 500 Boc Edwards) from Ge pellets (99.999%, Lesker) and highly ordered pyrolitic graphite (HOPG) sheets, respectively. Base vacuum was 5 10-7 mbar, which increased to 1 10-5 mbar during the evaporation processes. The aC/aGe/aC thin film was prepared without venting the chamber. The films were deposited onto freshly cleaved mica sheets and then transferred to Cu and SiN TEM grids in a water-based floating process for further inspection and in-situ device preparation.
2.3 In-situ device preparation
In order to study these materials, a piece of each has to be deposited between the two contact pads on the in-situ TEM chip, which was achieved with a dual-beam instrument (Thermo Fisher Scientific Helios 650) with SEM and focused ion beam (FIB). For bulk materials, FIB-based preparation methods for in-situ studies have been published [24] and their first step is the preparation of a conventional TEM lamella: a vertical slice of material with 1 m thickness, 12 m width and a depth of approximately 5 m is taken out from the bulk material and transferred to a TEM copper grid. In a second step, the slice is flipped and transferred to the contact pads of the in-situ chip, where a hole has been milled previously between the pads by FIB. The lamella can be thinned prior to transfer to the chip and/or directly on the chip by using a sample holder with inclined surface. Here, we did a combination of both. At first we thinned the sheet on the Cu grid down to 1 m thickness over the entire width and polished the surface that will be in contact to the chip to guarantee a good contact. Except for the CCO (see below) device, which was transferred without addtional pre-thinning, we further thinned the central part of the sheet from one side down to 700 nm forming a kind of bridge. This shape facilitates the thinning after the transfer of the flipped sheet to the chip as it moves the incidence of the focused ion beam away from the surface. The final thinning of the central part and the removal of the Pt cover from the TEM lamella preparation was done using an inclined sample holder that allows milling with the FIB parallel to the chips’ surface. SI2 describes this preparation exemplary for the CCO (see below) device.
For the MLC NT and the aGe thin-films, the preparation needs to be different and we employed a recently developed support-based method [22]. SI3 shows this preparation process exemplary for the MLC NT device, which consists in first preparing a conventional TEM sample by drop-casting the nanomaterial on a holey silicon nitride chip and the selection of a suitable NT by TEM. Using FIB, the NT and the support are transferred to the in-situ chip, where finally the silicon nitride is milled away to leave the NT as unique bridge between the contacts.
2.4 Electron microscopy
The MEMS chip is designed to fit in a DENSsolutions Wildfire holder with 4 pins for a ThermoFisherScientific/FEI instrument, which allows TEM analysis and electrical characterization of the chip at the same time. TEM, SAED and scanning (S)TEM analysis have been conducted in two aberration-corrected microscopes (Thermo Fisher Scientific Titan3 (image-corrected) and Titan low-base (probe-corrected)) operated at 300 keV electron energy. HAADF STEM imaging was done with a convergence and collection angle of 25 mrad and 48 mrad, respectively. Energy-dispersive X-ray spectroscopy (EDX) was performed in STEM mode with an Ultim X-MaxN 100TLE detector (Oxford Instruments). Electron energy-loss spectroscopy (EELS) data was acquired in STEM mode with a collection angle of 29.5 mrad using a Gatan Image Filter Tridiem 865 spectrometer.
During the electrical measurements, the specimen was irradiated with a relatively large parallel beam (TEM mode). An effect of the electron beam on the thermoelectric characterization could not be observed under these conditions. However, the device setup could allow to perform electron-beam induced current (EBIC) measurements or to investigate possible electron-beam heating effects.
2.5 Raman spectroscopy
A confocal Raman Alpha 300 M+ (WiTec) was used for acquisition of Raman spectra, with a 633 nm laser operated at 0.25 mW power and a 50x objective. The spectrometer was operated with 600 grooves/mm grating.
2.6 Electrical characterization
@@ 3.1 Reference materials and device characterization
Figure 2a shows a SEM image of the prepared p-doped Si device, where the material was cut into a cuboid with defined geometry in the central area above the hole in the membrane. The crystal orientation is 100 in vertical direction and approximately 032 in direction of the current flow between the contacts (SAED analysis in SI4). Energy-dispersive X-Ray spectroscopy (EDX) indicates a minor surface oxidation and Ga incorporation (SI5). Figure 2b depicts an I-V characterization of the device shown in (a) obtained at different applied heating currents between 0.1 (blue curve) and 5 mA (red curve).
With increasing applied to the differential heating device, the I-V curve shifts along the x-axis and the gradient slightly increases. The gradient change is linked to a slight decrease in average resistance from 890 to 775 due to an increase in temperature and a related improvement of the semiconducting Si and also likely due to a reduction of contact resistance. The shift can be attributed to the generation of a voltage induced by a temperature gradient along the sample. To measure this shift, a linear regression is performed on the I-V curves and the shift is calculated from and is marked by black circles in Figure 2b. The negative sign is due to the employed polarity (Figure 1c), as the temperature gradient is opposite to the applied voltage. The voltage offset is determined by the value of relative to (=0) to account for possible minor intrinsic currents. is shown in Figure 2c in dependence of for the Si device and a similar sample prepared from single-crystalline Mo. Both curves show a clear increase of the voltage offset with , which corresponds with the expected increase of the temperature gradient along the material. The voltage offset is positive, which agrees with the (positive) sign of the Seebeck coefficient of both materials [25, 26]. A quantitative assessment of the measured voltages is discussed below.
To check a possible contribution to the thermolectrical measurement by the chip itself, the raw in-situ chip, without deposited material, was characterized in the same way. No measurable contribution was observed (SI6).
@@ 3.2 Devices made of misfit-layered compounds
Figure 3 shows the structural analysis of the three devices made of misfit-layered compounds (MLCs). MLCs consist of two different layered materials stacked alternatively upon each other in c crystal direction, leading to a strong asymmetry between in-plane and out-of-plane directions of the crystal structure [1]. Two devices were prepared from the bulk CCO material with different crystal orientations with respect to the induced temperature gradient. Two SEM images of the first CCO device before and after FIB thinning for TEM studies are shown in Figure 3a. The microstructure of the Sr-doped CCO bulk material was analyzed using a standard TEM lamella (SI7), which revealed that the material possesses a micro-crystalline structure with most grains exhibiting a similar direction of the c-axis. The c-axis corresponds to the direction of the applied pressure during the hot-pressing process, see section 2.2. Grain sizes are small in c-direction (approximately 500 nm) and large in a-b directions (several m). A few grains and grain boundaries spanning horizontally across the device can be seen in the inset TEM image in Figure 3b, which also depicts a selected-area electron diffraction (SAED) pattern obtained from the thin part of the device. The SAED pattern confirms that the c-axis is perpendicular to the grain boundaries visible in Figure 3b, implying that for this device, the electrical current flow is predominantly in-plane and thus parallel to the layers of the MLC. This device is denoted as CCO in the following. The c-axis periodicity of the CCO material is measured to be 1.07 nm from the SAED pattern, which agrees with previously reported values on bulk CCO [27].
@@
Figure 3d-f details a second CCO device, denoted as CCO . Figure 3d shows a SEM image of the device, which was prepared such that the current flow is roughly perpendicular to the layered structure (parallel to the c-axis) and that a single grain boundary falls in the thinned central part. Figure 3d shows a bright-field (BF) TEM image of this device after orientation of the left grain into a high-order zone axis (TEM stage at =23°, =23°). The image clearly reveals the presence of two grains with the left grain oriented in zone-axis, appearing darker compared to the right grain. Figure 3e and 3f show two SAED patterns obtained from the left and right part of the device as marked in Figure 3d, corresponding to the two grains. The patterns confirm the orientation of the left grain in a higher-order zone-axis. An analysis of the reflections indicates the presence of [001] (1.07 nm) and [130] (0.15 nm) reflections, corresponding to a [-310] zone-axis orientation. Note that due to the monoclinic structure of the CCO material, [001] and [130] are not perpendicular to each other. On the other hand, the right grain does not exhibit a higher order orientation of the crystal with respect to the electron beam at this particular stage tilt (Figure 3f). The analysis of the device tilted to an angle where the right grain is oriented in [010] zone-axis (TEM stage at =16°, =19°) is shown in SI8. As the 001 direction roughly corresponds to the tilt direction, the angle between the c-axis directions of the two grains is estimated by the difference in tilt, which roughly corresponds to 7°. An EDX spectrum of this CCO device is shown in SI5 and confirms the presence of all elements of the CCO device, including the strontium doping.
@@
The third MLC device consists in a NT made of (LaS)1.14-TaS2. A SEM image of the device is shown in Figure 3c. The silicon nitride membrane utilized for the sample transfer to the chip was milled away in the central region in order that the NT forms the only conductive connection between the contact pads (see preparation in SI3). A rather long Pt-C focused ion beam induced deposition (FIBID) was needed to contact the NT to the upper contact pad. A SAED pattern acquired from the central part of the NT (Figure 3g) reveals that the complex structure of the MLC NT was preserved during the preparation [23]: Reflections are seen from the two different subsystems LaS (orange dashed rings) and TaS2 (blue dashed rings) as well as from the c-axis periodicity of the stack (pink arrow), which is perpendicular to the NT axis, as seen in Figure 3c. The tubular structure is recognized from the streaking reflections caused by the bending of the layers. The damage-free transfer with minimum Pt-C contamination is also seen from a high-angle annular dark field (HAADF) STEM image acquired from the edge of the NT (SI3) and an EDX spectrum of the MLC-NT device (SI5).